OpenStudy (anonymous):

Use synthetic substitution to evaluate the polynomial for the given value x = -3. P(x) = x3 + 3x2 + 4 P(-3) = ________ Enter a number only.

OpenStudy (campbell_st):

substitute x = -3 into the function P(-3) = (-3)^3 + 3(-3)^2 + 4 = -27 + 27 + 4 = 4 or by polynomial division  then Q(x) = x^2 R(x) = 4